# Statistical Inference Leveraging Synthetic Data with Distribution-Free Guarantees

### v1

@@ 3.3 GESPI for additional tasks: outlier detection and multiple testing
Multiple hypothesis testing. Another important statistical inference problem is multiple hypothesis testing, [^5] Multiple hypothesis testing has a broad range of applications across science and engineering, see for instance Benjamini and Hochberg (1995); Efron (2012); Bretz et al. (2016), etc. see e.g., Lehmann and Romano (2005b). Suppose we want to simultaneously test $m$ null hypotheses: $\mathcal{H}_{0,j}$, for $j=1,\ldots,m$. When testing many hypotheses at once—such as identifying outliers among a batch of test points—the probability of incorrectly labeling at least one inlier as an outlier can increase rapidly if each hypothesis is tested separately at level $\alpha$. This motivates the goal of simultaneously testing all $m$ nulls while controlling the family-wise error rate (FWER), or more generally, the $k$-FWER (Lehmann and Romano, 2005a) at level $\alpha$: $\mathbb{P}\left\{{\sum_{j=1}^{m}{\mathbbm{1}}\left\{{\text{$\mathcal{H}_{0,j}$ is true but rejected}}\right\}\geq k}\right\}\leq\alpha$, for some predetermined $k>0$.
@@ 3.4 The proposed GESPI framework
Now suppose we also have access to a synthetic/auxiliary dataset $\tilde{\mathcal{D}}_{N}=(\tilde{Z}_{1},\ldots,\tilde{Z}_{N})\in\mathcal{Z}^{N}$. Given a family of algorithms $\mathrm{Alg}_{\alpha}$ for each $\alpha\in(0,1)$ that attains the guarantee (5), we aim to construct an algorithm $\widetilde{\mathrm{Alg}}:\mathcal{Z}^{\infty}\times\mathcal{Z}^{\infty}\rightarrow\mathcal{A}$ that takes both $\mathcal{D}_{n}$ and $\tilde{\mathcal{D}}_{N}$ as input, such that the synthetic-leveraging procedure $\widetilde{\mathrm{Alg}}(\mathcal{D}_{n},\tilde{\mathcal{D}}_{N})$ improves upon the standard procedure $\mathrm{Alg}_{\cdot}(\mathcal{D}_{n})$.
3.4.1 General algorithm and theoretical guarantees
To introduce our method, we begin with a simpler setting than in our examples, where we only aim to upper bound the risk, and not to lower bound it; we will consider this setting below.
Condition 3.1 (informal).
@@ 3.4.1 General algorithm and theoretical guarantees
A formalized statement of Condition 3.1 is given in F.1. We note here that this is a mild condition, satisfied by all the inference problems discussed in this work; see Table˜3. For example, in hypothesis testing, the action space consists of reject/accept $\{0,1\}$ and the partial order $\preceq$ is defined by $\leq$. Recall that in the examples, our algorithm relies on taking unions and intersections or alternatively performing OR/AND operations. Generalizing our examples, our GESPI method takes the minimum ($\wedge$) of the output of two carefully chosen algorithms:
@@
where $\mathcal{D}_{n}\cup\tilde{\mathcal{D}}_{N}$ denotes the concatenated vector $(Z_{1},\ldots,Z_{n},\tilde{Z}_{1},\ldots,\tilde{Z}_{N})$.
Intuitively, the first component of the procedure (6), $\mathrm{Alg}_{\alpha}(\mathcal{D}_{n}\cup\tilde{\mathcal{D}}_{N})$, serves as the main part that incorporates the synthetic data, thereby producing a procedure based on a larger sample. The second component, $\mathrm{Alg}_{\alpha+\varepsilon}(\mathcal{D}_{n})$, at a relaxed level $\alpha+\varepsilon$, serves as a guardrail that does not depend on the synthetic data, and thus provides reliable statistical inference. The resulting procedure $\widetilde{\mathrm{Alg}}(\mathcal{D}_{n},\tilde{\mathcal{D}}_{N})$ tightly controls the risk when the synthetic distribution closely resembles the real one, while still guaranteeing risk control at level $\alpha+\varepsilon$ even when the synthetic data is of low quality.
@@
$\mathbb{E}_{{\mathcal{D}_{n}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P,\tilde{\mathcal{D}}_{N}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}Q,V\sim\mathcal{T}(P)}}\left[{\ell(\widetilde{\mathrm{Alg}}(\mathcal{D}_{n},\tilde{\mathcal{D}}_{N}),V)}\right]\leq\alpha+\min\{\varepsilon,c\cdot\textnormal{d}_{\ell,\mathrm{Alg}}(P,Q)\}\;\text{ for all }P,Q\in\mathcal{P},$
where  [^8] Here, $\textnormal{d}_{\textnormal{TV}}$ denotes the total variation distance. $\textnormal{d}_{\ell,\mathrm{Alg}}(P,Q)=\textnormal{d}_{\textnormal{TV}}(P_{\ell,\mathrm{Alg}}(P,Q),P_{\ell,\mathrm{Alg}}(Q,Q))$, and $P_{\ell,\mathrm{Alg}}(P,Q)$ denotes the distribution of $\ell(\mathrm{Alg}_{\alpha}(\mathcal{D}_{n}\cup\tilde{\mathcal{D}}_{N}),V)$ under $\mathcal{D}_{n}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$, $\tilde{\mathcal{D}}_{N}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}Q$, $\mathcal{V}\sim\mathcal{T}(P)$.
The above result provides a general upper bound, which depends on the quality of the synthetic data as measured by $\textnormal{d}_{\ell,\mathrm{Alg}}(P,Q)$. If the synthetic data are of high quality, this term is small, and the resulting risk is controlled close to the level $\alpha$. However, even if the synthetic data are of arbitrary poor quality, and $\textnormal{d}_{\ell,\mathrm{Alg}}(P,Q)\to\infty$, the guardrail is active, and the risk is controlled at $\alpha+\varepsilon$. Tighter and more interpretable bounds for specific applications are detailed in Appendix˜H.
Inference with two-sided guardrails. For the setting of two-sided guardrails, we have a similar version of GESPI, but one that also takes the maximum ($\vee$) with the output of the base algorithm:
@@
Suppose that Condition F.1 holds. Then given $\alpha,\varepsilon>0$, the algorithm $\widetilde{\mathrm{Alg}}$ defined in (7) deterministically satisfies $\mathrm{Alg}_{\alpha}(\mathcal{D}_{n})\preceq\widetilde{\mathrm{Alg}}(\mathcal{D}_{n},\tilde{\mathcal{D}}_{N})\preceq\mathrm{Alg}_{\alpha+\varepsilon}(\mathcal{D}_{n})$.
This result implies that the component $\mathrm{Alg}_{\alpha+\varepsilon}(\mathcal{D}_{n})$ serves as a guardrail for the validity of the synthetic-leveraged procedure $\smash{\widetilde{\mathrm{Alg}}(\mathcal{D}_{n},\tilde{\mathcal{D}}_{N})}$, while $\mathrm{Alg}_{\alpha}(\mathcal{D}_{n})$ ensures that $\smash{\widetilde{\mathrm{Alg}}(\mathcal{D}_{n},\tilde{\mathcal{D}}_{N})}$ is at least as useful as the base procedure $\mathrm{Alg}_{\alpha}(\mathcal{D}_{n})$—in terms of prediction interval width, test power, etc.
In the case of the hypothesis testing example (3), Theorem˜3.3 implies: $\smash{\widetilde{\mathrm{Alg}}(\mathcal{D}_{n},\tilde{\mathcal{D}}_{N})}\in\{0,1\}$ always rejects (=1=1) whenever $\mathrm{Alg}_{\alpha}(\mathcal{D}_{n})$ rejects (i.e., $\mathrm{Alg}_{\alpha}(\mathcal{D}_{n})\preceq\smash{\widetilde{\mathrm{Alg}}(\mathcal{D}_{n},\tilde{\mathcal{D}}_{N})}$), and never rejects unless $\mathrm{Alg}_{\alpha+\varepsilon}(\mathcal{D}_{n})$ rejects (i.e., $\smash{\widetilde{\mathrm{Alg}}(\mathcal{D}_{n},\tilde{\mathcal{D}}_{N})}\preceq\mathrm{Alg}_{\alpha+\varepsilon}(\mathcal{D}_{n})$). This implies that $\smash{\widetilde{\mathrm{Alg}}(\mathcal{D}_{n},\tilde{\mathcal{D}}_{N})}$’s Type I error is at most that of $\mathrm{Alg}_{\alpha+\varepsilon}(\mathcal{D}_{n})$ ($\leq\alpha+\varepsilon$), while its power is at least that of $\mathrm{Alg}_{\alpha}(\mathcal{D}_{n})$.
@@ 4 Experiments
We now demonstrate the performance of GESPI across several applications. Additional results are provided in Appendix˜C, which also include controlled experiments on simulated data to provide further insight into GESPI ’s performance. Full experimental details are provided in Appendix˜B.
@@
Methods. For each application, we compare the following methods: OnlyReal—the base inference method using only the real data. OnlySynth—the same inference method, but using only the synthetic data; this method does not have error rate control guarantees. GESPI—the proposed method that leverages both real and synthetic data, and is supported by error rate control guarantees.
@@ 4.1 Conformal risk control for protein structure prediction
We consider the protein structure prediction problem, where the input $X$ is the amino-acid (residue) sequence and the target $Y$ is the corresponding 3D structure (coordinates per residue). The goal is to control the proportion of residues whose prediction error exceeds a threshold (e.g., 3Å). We achieve this by abstaining predicted coordinates of residues that are likely to have such large prediction errors. Formally, we define a prediction set $C_{\lambda}(X)\subseteq X$ as the subset of the residues abstained on. We employ the conformal risk control framework (Angelopoulos et al., 2024; Bates et al., 2021) and utilize real data to tune a threshold $\hat{\lambda}$ such that $\mathbb{E}\left[{\frac{1}{|X|}\sum_{i\in X}\mathbb{I}\left\{\text{err}_{i}>3\text{\AA }\right\}\cdot\mathbb{I}\left\{i\notin C_{\hat{\lambda}}(X)\right\}}\right]\leq\alpha.$ Here, $\text{err}_{i}$ is the prediction error for residue $i$, which is formally defined in Section˜B.1. Note that the choice of 3Å is a standard scale for error, see, e.g., Jumper et al. (2021).
@@
Synthetic data. A key component of AlphaFold2 is the use of multi-sequence alignments (MSAs), where the model searches a terabyte-scale database for related sequences to improve predictions. Inspired by this, we treat the same MSAs used by AlphaFold2 as synthetic data ($\tilde{X}$) and generate corresponding predicted structures to approximate the prediction error, since true structures for the synthetic data are unavailable. The appeal of this construction is that we show how the powerful MSA component of AlphaFold2 can be utilized beyond its original purpose of improving predictions: we harness the MSA sequences to form high-quality synthetic data that boost statistical inference. Further details on the construction of the synthetic data are provided in Section˜B.1.
@@
We begin by visualizing the differences between the base OnlyReal method and our GESPI procedure. To illustrate how our method performs, we select protein T1029, for which the AlphaFold prediction is only partly accurate, and show the resulting prediction sets obtained by OnlyReal (Figure˜2(a)) and GESPI (Figure˜2(b)). Panel (1) shows the predicted structure, with residues abstained on highlighted in red and accepted residues in blue. Observe how OnlyReal conservatively abstains from all residues; this is a consequence of the limited real data used to tune $\lambda$. By contrast, GESPI abstains less, demonstrating the advantage of using synthetic data. Panel (2) shows the predicted structure (red and blue) aligned with the true structure (gray). Panel (3) highlights a small subset of residues for which the prediction is clearly inaccurate and where both methods abstain. Lastly, panel (4) shows a well-aligned region where OnlyReal abstains unnecessarily, while GESPI correctly accepts these residues. For completeness, we also visualize protein T1078 in Section˜C.2, where AlphaFold achieves relatively high accuracy.
(a) $\alpha=5\%$
(b) $\alpha=10\%$
Figure 3: Performance comparisons for protein structure prediction with error rate control. Conformal risk control methods applied at target levels (a) $\alpha=5\%$ and (b) $10\%$. Left: average risk (fraction of residues with error >3>3Å). Right: average abstention rate.
@@
Figure˜3 presents quantitative results for two $\alpha$ levels, showing a consistent trend: OnlyReal conservatively controls the risk but at the cost of a high abstention rate. In contrast, GESPI achieves risk close to the nominal $\alpha$ level while reducing the abstention rate. Crucially, OnlySynth serves only as a heuristic baseline and does not provide risk control guarantees. Additional results for $\alpha=15\%$ level, along with the selected thresholds for all $\alpha$ levels, are provided in Section˜C.2.
4.2 Hypothesis testing for win rate of large reasoning models
Given two large language models (LLMs), we aim to test whether model A outperforms model B on a specific type of task. Formally, we consider the hypotheses $\mathcal{H}_{0}:p=0.5\text{ vs. }\mathcal{H}_{1}:p>0.5,$ where $p$ denotes the win rate of model A over model B.
Data. We evaluate the win rate on the AIME25 dataset, which consists of 30 challenging reasoning math questions. The goal is to pin down the ranking of the models on AIME25 as closely as possible, which is hindered by the very small number of questions for this competition. For synthetic data, we use a subset of the OlympiadBench (He et al., 2024) math competition questions, which resemble AIME problems but are drawn from a different distribution and therefore cannot be treated as real test data. [^9] Since OlympiadBench was released in 2024, before AIME25, these two datasets are non-overlapping. For each question and model, we record whether the model’s answer is correct. Further details on the experimental setup are provided in Section˜B.2.
Figure 4: Performance comparisons for LLM win rate on AIME25 dataset. Hypothesis testing methods applied at level $\alpha=5\%$ and $\varepsilon=2\%$. Left: Description of model comparisons. Middle: Power, comparing the rejection rate under the standard setting. Right: Type I error, measured in the shuffled-response setting where the null holds.
Figure˜4 presents two model comparisons, each involving a distinct pair of LLMs. In both cases, the rejection rate (power) is well above the target level $\alpha$, indicating that the first model outperforms the second. Our proposed method, GESPI, achieves higher power than OnlyReal, while in the shuffled-answers setting, both methods achieve Type I error at the target level $\alpha$.
4.3 Outlier detection with contaminated reference set

### v2

@@ 3.3 GESPI for Multiple Hypothesis Testing
Suppose we want to simultaneously test $m$ null hypotheses: $\mathcal{H}_{0,j}$, for $j=1,\ldots,m$. When testing many hypotheses at once—such as identifying outliers among a batch of test points—the probability of incorrectly labeling at least one inlier as an outlier can increase rapidly if each hypothesis is tested separately at level $\alpha$. This motivates the goal of simultaneously testing all $m$ nulls while controlling the family-wise error rate (FWER), or more generally, the $k$-FWER (Lehmann and Romano, 2005a) at level $\alpha$: $\mathbb{P}\left\{{\sum_{j=1}^{m}{\mathbbm{1}}\left\{{\text{$\mathcal{H}_{0,j}$ is true but rejected}}\right\}\geq k}\right\}\leq\alpha$, for some predetermined $k>0$.
@@ 3.4 The Proposed GESPI Framework
Now suppose we also have access to a synthetic/auxiliary dataset $\tilde{\mathcal{D}}_{N}=(\tilde{Z}_{1},\ldots,\tilde{Z}_{N})\in\mathcal{Z}^{N}$. Given a family of algorithms $\mathrm{Alg}_{\alpha}$ for each $\alpha>0$ that attains the guarantee (4), we aim to construct an algorithm $\widetilde{\mathrm{Alg}}:\mathcal{Z}^{\infty}\times\mathcal{Z}^{\infty}\rightarrow\mathcal{A}$ that takes both $\mathcal{D}_{n}$ and $\tilde{\mathcal{D}}_{N}$ as input, such that the synthetic-leveraging procedure $\widetilde{\mathrm{Alg}}(\mathcal{D}_{n},\tilde{\mathcal{D}}_{N})$ improves upon the standard procedure $\mathrm{Alg}_{\cdot}(\mathcal{D}_{n})$.
3.4.1 General Algorithm and Theory
To introduce our method, we begin with a simpler procedure, which only aim to upper bound the risk, and not to lower bound it; later we will consider the formulation presented in the examples.
Condition 3.2 (informal).
@@ 3.4.1 General Algorithm and Theory
A formalized statement of Condition 3.2 is given in F.1. Intuitively, this condition allows us to robustly aggregate real and synthetic-powered actions by (i) comparing them, and (ii) choosing the better action in the view of the risk we aim to control. We note here that this is a mild condition, satisfied by all the inference problems discussed in this work; see Table˜S2. For example, in hypothesis testing, the action space consists of accept/reject $\{0,1\}$ and the partial order $\preceq$ is defined by $\leq$. Recall that in the examples, our algorithm relies on taking unions and intersections or alternatively performing OR/AND operations. Generalizing our examples, our GESPI method takes the minimum ($\wedge$) of the output of two carefully chosen algorithms:
@@
where $\mathcal{D}_{n}\cup\tilde{\mathcal{D}}_{N}$ is the vector $(Z_{1},\ldots,Z_{n},\tilde{Z}_{1},\ldots,\tilde{Z}_{N})$.
Intuitively, the first component of (5), $\mathrm{Alg}_{\alpha}(\mathcal{D}_{n}\cup\tilde{\mathcal{D}}_{N})$, serves as the main part that incorporates the synthetic data, thereby producing a procedure based on a larger sample. The second component, $\mathrm{Alg}_{\alpha+\varepsilon}(\mathcal{D}_{n})$, at a relaxed level $\alpha+\varepsilon$, serves as a guardrail that does not depend on the synthetic data, and thus provides reliable statistical inference. The resulting procedure $\widetilde{\mathrm{Alg}}(\mathcal{D}_{n},\tilde{\mathcal{D}}_{N})$ tightly controls the risk when the synthetic distribution closely resembles the real one, while still guaranteeing risk control at level $\alpha+\varepsilon$ even when the synthetic data are of low quality.
@@
$\mathbb{E}_{{\mathcal{D}_{n}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P,\tilde{\mathcal{D}}_{N}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}Q,V\sim\mathcal{T}(P)}}\left[{\ell(\widetilde{\mathrm{Alg}}(\mathcal{D}_{n},\tilde{\mathcal{D}}_{N}),V)}\right]\leq\alpha+\min\{\varepsilon,c\cdot\textnormal{d}_{\ell,\mathrm{Alg}}(P,Q)\}$
for all $P,Q\in\mathcal{P}$, where [^8] Here, $\textnormal{d}_{\textnormal{TV}}$ denotes the total variation distance. $\textnormal{d}_{\ell,\mathrm{Alg}}(P,Q)=\textnormal{d}_{\textnormal{TV}}(P_{\ell,\mathrm{Alg}}(P,Q),P_{\ell,\mathrm{Alg}}(Q,Q))$, and $P_{\ell,\mathrm{Alg}}(P,Q)$ denotes the distribution of $\ell(\mathrm{Alg}_{\alpha}(\mathcal{D}_{n}\cup\tilde{\mathcal{D}}_{N}),V)$ under $\mathcal{D}_{n}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$, $\tilde{\mathcal{D}}_{N}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}Q$, $\mathcal{V}\sim\mathcal{T}(P)$.
The above result provides a general upper bound that depends on the synthetic data quality, as measured by $\textnormal{d}_{\ell,\mathrm{Alg}}(P,Q)$. If the synthetic data are of high quality, this term is small and the risk is controlled close to level $\alpha$. However, even if the synthetic data are of arbitrary poor quality and $\textnormal{d}_{\ell,\mathrm{Alg}}(P,Q)\to\infty$, the guardrail ensures risk control at $\alpha+\varepsilon$. Tighter and more interpretable bounds for specific applications are detailed in Appendix˜H.
Inference with two-sided guardrails. For this setting, we have a similar version of GESPI, but one that also takes the maximum ($\vee$) with the output of the base algorithm:
@@
Suppose that Condition F.1 holds. Then given $\alpha,\varepsilon>0$, the algorithm defined in (6) deterministically satisfies $\mathrm{Alg}_{\alpha}(\mathcal{D}_{n})\preceq\widetilde{\mathrm{Alg}}(\mathcal{D}_{n},\tilde{\mathcal{D}}_{N})\preceq\mathrm{Alg}_{\alpha+\varepsilon}(\mathcal{D}_{n})$.
This result implies that the component $\mathrm{Alg}_{\alpha+\varepsilon}(\mathcal{D}_{n})$ serves as a guardrail for the validity of the synthetic-leveraged procedure $\smash{\widetilde{\mathrm{Alg}}(\mathcal{D}_{n},\tilde{\mathcal{D}}_{N})}$, while $\mathrm{Alg}_{\alpha}(\mathcal{D}_{n})$ ensures that $\smash{\widetilde{\mathrm{Alg}}(\mathcal{D}_{n},\tilde{\mathcal{D}}_{N})}$’s utility is at least that of the base procedure $\mathrm{Alg}_{\alpha}(\mathcal{D}_{n})$ (e.g., prediction interval width, test power, etc).
@@ 4 Experiments
We now demonstrate the performance of GESPI across several applications. Additional results and applications—including comparison of LLMs and mechanistic interpretability of a Vision Transformer model—are provided in Appendix˜C. This section also includes controlled experiments on simulated data to provide further insight into GESPI ’s performance. Full experimental details are provided in Appendix˜B.
@@
Methods. For each application, we compare the following methods: OnlyReal—the base inference method using only real data; Guardrail—same as OnlyReal, but with a higher error level $\alpha+\varepsilon$; OnlySynth—the same inference method using only synthetic data; this method lacks error rate control guarantees and is included solely to illustrate the unknown quality of the synthetic data; GESPI—the proposed method that leverages both real and synthetic data, and supported by error rate control guarantees.
In Appendix˜J we include an additional baseline—denoted Synth+Real—that pools the real and synthetic data and then applies the base inference method at level $\alpha$. This baseline does not have any error rate control guarantees, similarly to OnlySynth, and therefore does not constitute a competitive method; it is included only to illustrate the unknown quality of the pooled data. Moreover, in the regime $N\gg n$—which is our main setting—OnlySynth and Synth+Real perform similarly.
4.1 Conformal Prediction for Image Classification
We begin with an image classification task, where the goal is to obtain class-conditional coverage. Specifically, given a test image $X_{n+1}$, we aim to construct a prediction set $\hat{C}(X_{n+1})$ such that
$\mathbb{P}\left\{{Y_{n+1}\in\hat{C}(X_{n+1})\mid Y_{n+1}=y}\right\}\geq 1-\alpha,\text{for each }y\in\mathcal{Y}.$
The purpose of this section is to better understand and visualize how GESPI adapts to the quality of the synthetic data, and how this quality affects the performance compared to using only the real data.
Data and experimental setup. We follow the setup introduced in Bashari et al. (2025a). We use a subset of 30 classes from ImageNet as the real data and their corresponding FLUX-generated images as the synthetic data Black Forest Labs (2024). The target coverage level is $1-\alpha=95\%$ with $\varepsilon=2\%$. Additional details are in Section˜B.1.
For this particular setting of the hyperparameters, SPI Bashari et al. (2025a) performs similarly to our method, as $N\gg n$; so we omit the results. Note, however, that this is the only application among those considered in this paper where SPI can be applied. That said, we include a comparison with SPI for different values of $N$ in Section˜C.2. There, we show that for smaller values of $N$, GESPI outperforms SPI. This is because our synthetic-powered prediction sets are constructed using the pooled data and not merely using the synthetic data as in SPI.
Figure˜3 presents the performance of various conformal methods for selected classes. OnlyReal conservatively controls the coverage at level $1-\alpha$, producing trivial, non-informative prediction sets due to the limited sample size. In contrast, Guardrail controls the coverage at level $1-\alpha-\varepsilon$, though it exhibits high variability for the same reason. OnlySynth obtains arbitrary coverage levels with low variability, depending on the (unknown) synthetic data quality. The deviation of the observed coverage from the nominal level $1-\alpha$ serves as a proxy for synthetic data quality: the smaller the deviation, the higher the quality.
In turn, GESPI controls the coverage in the worst-case at level $1-\alpha-\varepsilon$, while achieving coverage close to the nominal level $1-\alpha$ with low variability when the synthetic data are of high quality. By design, GESPI handles the full spectrum of synthetic data quality—from worst to best—without requiring any a-priori knowledge of that quality.
Figure 3: Performance comparisons for image classification with class-conditional coverage on ImageNet. Conformal prediction methods applied at level $\alpha=5\%$ and $\varepsilon=2\%$. FLUX-generated images serve as the synthetic data. Results are shown for selected classes; see Tables˜S3 and S4 for all classes.
This adaptivity allows GESPI to achieve coverage close to the nominal level with low variability, rather than always attaining $1-\alpha-\varepsilon$ coverage as Guardrail does. The variance reduction is practically important, as practitioners typically have access to only a single dataset. A lower variance ensures that the holdout-data-conditional coverage is more likely to be close to the target level $1-\alpha$.
For example, for the Magpie class, the variance of Guardrail is extremely large due to the small sample size, resulting in realized coverage as low as 60%, despite the target level being 95%. In contrast, GESPI achieves coverage close to 95% with substantially lower variability.
To further illustrate that the quality of synthetic data is non-trivial to assess—and specifically, unknown in practice—Figure˜S6 shows example FLUX-generated images for the Siberian Husky class. Interestingly, although these synthetic images appear to be of a high quality, they do not prove useful for our current task; further emphasizing the importance of safely leveraging synthetic data. Indeed, as shown in Figure˜3, OnlySynth achieves only 60% coverage, well below the target level of $95\%$, indicating that the synthetic data are of low quality.
We include additional results in Section˜C.2, including a comparison of GESPI and OnlyReal as a function of empirical coverage. In this experiment, we artificially align the two approaches at the same mean empirical coverage to isolate the effect of using additional high-quality synthetic data. This illustrate the variance reduction benefit of GESPI compared to using only real data. Of course, such a comparison is only illustrative, as it requires knowing the empirical coverage, which is not feasible in practice.
@@ 4.2 Conformal Risk Control for Protein Structure Prediction
We consider the protein structure prediction problem, where the input $X$ is the amino-acid (residue) sequence and the target $Y$ is the corresponding 3D structure (coordinates per residue). The goal is to control the proportion of residues whose prediction error exceeds a threshold (e.g., 3Å). We achieve this by abstaining predicted coordinates of residues that are likely to have such large prediction errors. Formally, we define a prediction set $C_{\lambda}(X)\subseteq X$ as the subset of the residues abstained on. We employ the conformal risk control framework (Angelopoulos et al., 2024; Bates et al., 2021) and utilize real data to tune a threshold $\hat{\lambda}$ such that $\mathbb{E}\left[{\frac{1}{|X|}\sum_{i\in X}\mathbb{I}\left\{\text{err}_{i}>3\text{\AA }\right\}\cdot\mathbb{I}\left\{i\notin C_{\hat{\lambda}}(X)\right\}}\right]\leq\alpha.$ Here, $\text{err}_{i}$ is the prediction error for residue $i$, which is formally defined in Section˜B.2. Note that the choice of 3Å is a standard scale for error, see, e.g., Jumper et al. (2021).
@@
Synthetic data. A key component of AlphaFold2 is the use of multi-sequence alignments (MSAs), where the model searches a terabyte-scale database for related sequences to improve predictions. Inspired by this, we treat the same MSAs as synthetic data ($\tilde{X}$) and generate predicted structures to approximate the prediction error, since true structures for the synthetic data are unavailable. The appeal of this construction is that we show how the powerful MSA component of AlphaFold2 can be utilized beyond its original purpose of improving predictions: we harness the MSA sequences to form high-quality synthetic data that boost statistical inference. Further details on the construction of the synthetic data are provided in Section˜B.2.
@@
We begin by visualizing the differences between the base OnlyReal method and our GESPI procedure. To illustrate how our method performs, we select protein T1029, where AlphaFold’s prediction is only partly accurate, and show the resulting prediction sets obtained by OnlyReal (Figure˜2(a)) and GESPI (Figure˜2(b)). Panel (1) shows the predicted structure, with residues abstained on highlighted in red and accepted residues in blue. Observe how OnlyReal conservatively abstains from all residues; this is a consequence of the limited real data used to tune $\lambda$. By contrast, GESPI abstains less, demonstrating the advantage of using synthetic data. Panel (2) shows the predicted structure (red and blue) aligned with the true structure (gray). Panel (3) highlights a small subset of residues for which the prediction is clearly inaccurate and where both methods abstain. Lastly, panel (4) shows a well-aligned region where OnlyReal abstains unnecessarily, while GESPI correctly accepts these residues. For completeness, we also visualize protein T1078 in Section˜C.3, where AlphaFold achieves relatively high accuracy.
@@
Figure˜4 presents quantitative results for two $\alpha$ levels, showing a consistent trend: OnlyReal conservatively controls the risk but at the cost of a high abstention rate. Guardrail controls the risk at a higher level $\alpha+\varepsilon$ but exhibits high variability due to limited real data. In contrast, GESPI achieves risk close to the nominal $\alpha$ level with lower variability, while reducing the abstention rate compared to OnlyReal. Crucially, OnlySynth serves only as a heuristic baseline and does not provide risk control guarantees. Additional results for $\alpha=10\%$, along with the selected thresholds for all $\alpha$ levels, are in Section˜C.3.
Figure 4: Performance comparisons for protein structure prediction with error rate control. Conformal risk control methods at $\alpha=5\%$ and $15\%$. Left: average risk (fraction of residues with error >3>3Å). Right: average abstention rate.
4.3 Multiple Hypothesis Testing for Outlier Detection